OpenStudy (anonymous):

Simplify 6(2r − 5) + 2r

OpenStudy (anonymous):

Step 1: Multiply multiply the 6, 6(2r - 5) + 2r = (6 x 2r - 6 x 5) + 2r = (12r - 30) + 2r Step 2: Combine like terms = 12r - 30 + 2r = 12r +2r - 30 = 14r - 30 Solution: 14r - 30 Hope that helps

OpenStudy (anonymous):

thankyou helps alot !:)

OpenStudy (anonymous):

your welcome glad i could help :)